# Chow polynomials of totally nonnegative matrices and posets

### v1

@@ 4.3. Chow-deranged maps and Chow-Eulerian transformations
For $R_{0}=(\binom{n}{k})_{n,k=0}^{\infty}$, the deranged map was first studied by the first author and Solus [BS21], who proved Theorem 4.14 below for the case when $R=R_{0}$. The Chow-Eulerian map for the case when $R=R_{0}$ was first considered by Brenti [Bre89], who conjectured that $\mathcal{A}_{R_{0}}(f)$ is real-rooted whenever all zeros of $f\in\mathbb{R}[t]$ are real and nonpositive. This conjectured was disproved in [BJ22], where it was conjectured that $\mathcal{A}_{R_{0}}(f)$ is real-rooted whenever $f$ has a nonnegative expansion in $\{t^{k}(1+t)^{n-k}\}_{k=0}^{n}$. This conjecture was proved by Athanasiadis in [Ath23]. Theorem 4.15 below generalizes Athanasiadis result to any resolvable matrix. Suppose $R$ is resolvable, and define
@@
The proof is by induction over $n$, the case when $n=0$ being clear. By induction the sequence $\{\lambda_{n,j}d_{n,j}\}_{j=0}^{n}$ is $\mathcal{I}_{n}$-interlacing. The proof now follows by Lemma 4.13, Theorem 4.12 and induction. ∎
@@ 5.1. Coefficients of $\gamma$-Chow polynomials of matrices
Given a lower triangular matrix $R=(r_{n,k})_{n,k=0}^{N}$ with all diagonal entries equal to one, we know that the Chow polynomials $H_{n}=H_{n}[R]$ are palindromic with center of symmetry (n1)/2(n-1)/2. We write $\gamma_{n}[R]=\gamma(H_{n}[R])$ for the $\gamma$-polynomial of $H_{n}$ and we call them the $\gamma$-Chow polynomials of the matrix $R$. The main goal of this section is to give an interpretation to the coefficients of these polynomials. For a set $S=\{s_{1}<\cdots<s_{k}\}\subseteq[n-1]$} define
@@
A set $S$ of integers is called stable if $S$ does not contain any two consecutive integers.
@@
$\gamma_{n}[R]$ $=\sum_{\begin{subarray}{c}T=\{s_{1}<\cdots<s_{m}\}\subseteq[n-1]\\ T\cup\{0\}\text{ stable}\end{subarray}}(-1)^{|T|}(-t)^{|T|}\alpha_{[0,n]}(T)\cdot\left[W_{n-j_{m}-1}\ \prod_{i=1}^{m}W_{j_{i}-j_{i-1}-1}\right]$
@@
where in the second equality we used [FMV24, Lemma 4.23]. In the third equality we exchanged the two sums and used that if $S\cup\{0\}$ is stable then all of its subsets are also stable. ∎
@@
5.2. Interlacing $\gamma$-Chow polynomials
Next we will prove some refined interlacing results of Chow polynomials of $\operatorname{\mathrm{TN}}$-matrices. These will be used in Section 7 to prove that Chow polynomials of paving matroids are real-rooted.
@@ 6. Chow polynomials of $\mathrm{TN}$-posets
  • Boolean cell complex with nonnegative $h$-vectors. These include Cohen-Macaulay simplicial complexes and face lattices of simplicial polytopes.
@@
  • $q$-poset with nonnegative $h$-vectors [Ald10]. These include shellable $q$-complexes [GPR22].
  • Perfect matroid designs [Dez92], i.e., rank uniform geometric lattices. These include projective geometries and affine geometries.
@@
Recall that the characteristic augmented Chow polynomial of a poset $P$ is equal to the one of its dual $P^{*}$ by [BV25, Corollary 4.2]. Then one has the following straightforward consequence.
@@ 7. Chow polynomials of lattices of flats of paving matroids
8. Toeplitz matrices and Pólya frequency sequences
In this section we will focus on the special case when $R$ is a Toeplitz matrix. Let $\{a_{n}\}_{n=0}^{\infty}$ be a sequence of elements in an integral domain $\mathrm{R}$, where $a_{0}=1$. Consider the lower triangular Toeplitz matrix $R=(a_{n-k})_{n,k=0}^{\infty}$, where $a_{m}=0$ if $m<0$.
Consider the formal power series $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ in $\mathrm{R}[[z]]$, and define formal power series in $\mathrm{R}[t][[z]]$ by
@@ 8. Toeplitz matrices and Pólya frequency sequences
Let $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ be a formal power series in $\mathrm{R}[[z]]$, where $a_{0}=1$. Then
@@
Let $R$ be the Toeplitz matrix corresponding to $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$. For $n\geq 0$ and $k\geq 1$, let $R(n,k)$ denote the matrix obtained from $R$ by deleting all rows and columns indexed by $n,n+1,\ldots,n+k-1$. Then
@@
$f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}=Cz^{N}e^{\gamma z}\prod_{i=1}^{\infty}\frac{1+\alpha_{i}z}{1-\beta_{i}z},$ (8.1)
@@
Applying Theorem 4.16 to Toeplitz matrices produces four families of real-rooted polynomials to any series $f(z)$ of the form (8.1).
@@
Suppose $f(z)$ is a power series as in (8.1). Then the polynomials $d_{n}(t),H_{n}(t),A_{n}(t)$ and $G_{n}(t)$, $n\geq 0$, defined via $f(z)$ by the identities in Theorem 8.1 are all real-rooted.
@@
When $a_{0}>0$, then the theorem follows immediately from Theorems 8.3 and 4.16. The case when $a_{0}=0$, follows by continuity and Hurwitz’ theorem on the continuity of zeros [Sch05, Theorem 1.3.8] by considering
$f_{\epsilon}(z)=C(\epsilon+z)^{N}e^{\gamma z}\prod_{i=1}^{\infty}\frac{1+\alpha_{i}z}{1-\beta_{i}z},$
@@
8.1. Binomial posets
Recall [Sta76, Sta12] that a binomial poset is a locally finite poset $P$ for which
@@ 8.1. Binomial posets
$\sum_{n=0}^{\infty}d_{n}(t)\frac{z^{n}}{B(n)}$ $=\frac{1-t}{f(tz)-tf(z)},\quad\quad\sum_{n=0}^{\infty}H_{n}(t)\frac{z^{n}}{B(n)}$ $=\frac{(1-t)f(z)}{f(tz)-tf(z)},$
$\sum_{n=0}^{\infty}A_{n}(t)\frac{z^{n}}{B(n)}$ $=\frac{(1-t)f(tz)}{f(tz)-tf(z)},\quad\quad\sum_{n=0}^{\infty}G_{n}(t)\frac{z^{n}}{B(n)}$ $=\frac{(1-t)f(tz)f(z)}{f(tz)-tf(z)}.$
@@
Consider $\mathbb{B}(q)$, the lattice of all finite dimensional subspaces of an infinite dimensional vector space over a field with $q$ elements. Then $\mathbb{B}(q)$ is boolean with factorial function given by
@@ 8.2. Symmetric and supersymmetric functions
Let $\mu\subseteq\lambda$ be partitions. Then the supersymmetric skew Schur function, $s_{\lambda/\mu}(\mathbf{x}/\mathbf{y})$, may be defined by the Jacobi-Trudi identity as
@@
We say that $f(\mathbf{x}/\mathbf{y};t)$ is super Schur $\gamma$-positive if each $\gamma_{k}(\mathbf{x}/\mathbf{y})$ has a nonnegative expansion in super Schur functions. Notice that $f(\mathbf{x}/\mathbf{y};t)\in\Lambda(\mathbf{x}/\mathbf{y})$ is super Schur $\gamma$-positive if and only if $f(\mathbf{x}/0;t)$ is Schur $\gamma$-positive if and only if $f(/\mathbf{y};t)$ is Schur $\gamma$-positive. This is because the maps from supersymmetric functions to symmetric functions defined by
@@
$s_{\lambda/\mu}(\mathbf{x}/\mathbf{y})=\sum_{\nu}c_{\mu\nu}^{\lambda}s_{\nu}(\mathbf{x}/\mathbf{y}),$

### v2

@@ 4.3. Chow-deranged maps and Chow-Eulerian transformations
Remark 4.14.
For $R_{0}=(\binom{n}{k})_{n,k=0}^{\infty}$, the deranged map was first studied by the first author and Solus [BS21], who proved Theorem 4.16 below for the case when $R=R_{0}$. The Chow-Eulerian map for the case when $R=R_{0}$ was first considered by Brenti [Bre89], who conjectured that $\mathcal{A}_{R_{0}}(f)$ is real-rooted whenever all zeros of $f\in\mathbb{R}[t]$ are real and nonpositive. This conjectured was disproved in [BJ22], where it was conjectured that $\mathcal{A}_{R_{0}}(f)$ is real-rooted whenever $f$ has a nonnegative expansion in $\{t^{k}(1+t)^{n-k}\}_{k=0}^{n}$. This conjecture was proved by Athanasiadis in [Ath23].
Theorem 4.17 below generalizes Athanasiadis result to any resolvable matrix. Suppose $R$ is resolvable, and define
@@
Moreover if $f=\sum_{k=0}^{n}h_{k}R_{n,k}$, where $h_{k}\geq 0$ for each $k$, then $\mathcal{D}(f)$ is real-rooted and $d_{n,0}\prec\mathcal{D}(f)\prec d_{n,n}$.
@@
The proof of the first statement is by induction over $n$, the case when $n=0$ being clear. By induction the sequence $\{\lambda_{n,j}d_{n,j}\}_{j=0}^{n}$ is $\mathcal{I}_{n}$-interlacing. The proof of the first statement now follows by Lemma 4.15, Theorem 4.13 and induction.
The proof of the second statement from the first in conjunction with Proposition 4.7. ∎
@@
Moreover if $f=\sum_{k=0}^{n}h_{k}R_{n,k}$, where $h_{k}\geq 0$ for each $k$, then $\mathcal{A}(f)$ is real-rooted and $A_{n,0}\prec\mathcal{A}(f)\prec A_{n,n}$.
@@ 5.1. Coefficients of $\gamma$-Chow polynomials of matrices
Given a lower triangular matrix $R=(r_{n,k})_{n,k=0}^{N}$ with all diagonal entries equal to one, we know that the Chow polynomials $H_{n}=H_{n}[R]$ are palindromic with center of symmetry (n1)/2(n-1)/2. We write $\gamma_{n}[R]=\gamma(H_{n}[R])$ for the $\gamma$-polynomial of $H_{n}[R]$, and we call it a $\gamma$-Chow polynomials of the matrix $R$. The main goal of this section is to give an interpretation to the coefficients of these polynomials. For a set $S=\{s_{1}<\cdots<s_{k}\}\subseteq[n-1]$} define
@@
A set $S$ of integers is called stable if it does not contain any two consecutive integers.
@@
$\gamma_{n}[R]$ $=\sum_{\begin{subarray}{c}T=\{s_{1}<\cdots<s_{m}\}\subseteq[n-1]\\ T\cup\{0\}\text{ stable}\end{subarray}}(-1)^{|T|}(-t)^{|T|}\alpha_{[0,n]}(T)\cdot\left(W_{n-j_{m}-1}\ \prod_{i=1}^{m}W_{j_{i}-j_{i-1}-1}\right)$
@@
where in the second equality we used [FMV24, Lemma 4.23]. In the third equality we exchanged the two sums and used that if $S\cup\{0\}$ is stable, then all of its subsets are also stable. ∎
@@
5.2. Interlacing properties of $\gamma$-Chow polynomials
Next we will prove some refined interlacing properties satisfied by $\gamma$-Chow polynomials of $\operatorname{\mathrm{TN}}$-matrices. These will be used in Section 7 to prove that Chow polynomials of paving matroids are real-rooted.
@@ 6. Chow polynomials of $\mathrm{TN}$-posets
  • a.
    Boolean cell complex with nonnegative $h$-vectors. For example Cohen-Macaulay simplicial complexes and face lattices of simplicial polytopes.
  • b.
@@
  • c.
    $q$-poset with nonnegative $h$-vectors [Ald10]. For example shellable $q$-complexes [GPR22].
  • d.
    Perfect matroid designs [Dez92], i.e., rank uniform geometric lattices. These include (truncations of) projective geometries and affine geometries.
  • e.
@@
If we apply Theorem 6.1 to a above, then we recover the result of Hoster and Stump [HS25].
@@
Recall that the characteristic augmented Chow polynomial of a poset $P$ is equal to the one of its dual $P^{*}$ by [BV25, Corollary 4.2], from which the following consequence follows.
@@ 7. Chow polynomials of lattices of flats of paving matroids
8. Toeplitz matrices
In this section we will focus on the special case when $R$ is a Toeplitz matrix. Let $\{a_{n}\}_{n=0}^{\infty}$ be a sequence of elements in an integral domain $\mathrm{R}$, where $a_{0}=1$. Associate to $\{a_{n}\}_{n=0}^{\infty}$ the lower triangular Toeplitz matrix $R=(a_{n-k})_{n,k=0}^{\infty}$, where $a_{m}=0$ if $m<0$.
Consider the formal power series $f=\sum_{n=0}^{\infty}a_{n}z^{n}$ in $\mathrm{R}[[z]]$, and define formal power series in $\mathrm{R}[t][[z]]$ by
@@ 8. Toeplitz matrices
Let $f=\sum_{n=0}^{\infty}a_{n}z^{n}$ be a formal power series in $\mathrm{R}[[z]]$, where $a_{0}=1$. Then
@@
Let $R$ be the Toeplitz matrix corresponding to $f=\sum_{n=0}^{\infty}a_{n}z^{n}$. For $n\geq 0$ and $k\geq 1$, let $R(n,k)$ denote the matrix obtained from $R$ by deleting all rows and columns indexed by $n,n+1,\ldots,n+k-1$. Then
@@
$f=\sum_{n=0}^{\infty}a_{n}z^{n}=Cz^{N}e^{\gamma z}\prod_{i=1}^{\infty}\frac{1+\alpha_{i}z}{1-\beta_{i}z},$ (8.1)
@@
Applying Theorem 4.18 to Toeplitz matrices produces four families of real-rooted polynomials to any series $f$ of the form (8.1).
@@
Suppose $f$ is a power series as in (8.1). Then the polynomials $d_{n}(t),H_{n}(t),A_{n}(t)$ and $G_{n}(t)$, $n\geq 0$, defined via $f$ by the identities in Theorem 8.1 are all real-rooted.
@@
When $a_{0}>0$, then the theorem follows immediately from Theorems 8.3 and 4.18. The case when $a_{0}=0$, follows by continuity and Hurwitz’ theorem on the continuity of zeros [Sch05, Theorem 1.3.8] by considering
$f_{\epsilon}=C(\epsilon+z)^{N}e^{\gamma z}\prod_{i=1}^{\infty}\frac{1+\alpha_{i}z}{1-\beta_{i}z},$
@@
8.1. Binomial and Sheffer posets
Recall [DRS72, Sta76, Sta12] that a binomial poset is a locally finite poset $P$ for which
@@ 8.1. Binomial and Sheffer posets
$\sum_{n=0}^{\infty}d_{n}(t)\frac{z^{n}}{B(n)}$ $=\frac{1-t}{b(tz)-tb(z)},\quad\quad\sum_{n=0}^{\infty}H_{n}(t)\frac{z^{n}}{B(n)}$ $=\frac{(1-t)b(z)}{b(tz)-tb(z)},\qquad$
$\sum_{n=0}^{\infty}A_{n}(t)\frac{z^{n}}{B(n)}$ $=\frac{(1-t)b(tz)}{b(tz)-tb(z)},\quad\quad\sum_{n=0}^{\infty}G_{n}(t)\frac{z^{n}}{B(n)}$ $=\frac{(1-t)b(tz)\cdot b(z)}{b(tz)-tb(z)}.$
@@
Let $\mathbb{F}_{q}$ be a field with $q$ elements, and let $V(q)$ be the free $\mathbb{F}_{q}$-linear space over the set $\{e_{1},e_{2},\ldots\}$. Let further $\mathbb{B}(q)$ be the lattice of all finite dimensional subspaces of $V(q)$. Then $\mathbb{B}(q)$ is boolean with factorial function given by
@@
Ehrenborg and Readdy generalized the notion of binomial posets in [ER95]. A poset $P$ is called a Sheffer poset if there are two functions $B:\mathbb{N}\to\mathbb{N}$ and $C:\mathbb{N}\to\mathbb{N}$ such that
  1. (1)
    $P$ is locally finite with a least element $\hat{0}$, and $P$ contains an infinite chain.
  2. (2)
    Each interval in $P$ is graded, and hence $P$ has a rank function $\rho:P\to\mathbb{N}$.
  3. (3)
    The number of maximal chains in $\hat{0},x]$ is equal to $C(\rho(x))$, for each $x\in P$.
  4. (4)
    If $\hat{0}<x\leq y$, then the number of maximal chains in $[x,y]$ is equal to $B(\rho(x,y))$.
The functions $B$ and $C$ are called the factorial functions associated to $P$. Hence Sheffer posets are rank uniform, and binomial posets are the Sheffer posets for which $B=C$.
Example 8.7.
The following posets are Sheffer. For more example, see [ER95].
  • a.
    Let $r$ be a positive integer. The infinite $r$-cubical lattice is the infinite direct product (with a least element $\hat{0}$ adjoined)
    $\mathbf{C}_{r}=\{\hat{0}\}\cup\prod_{n=1}^{\infty}N_{r},$
    where $N_{r}$ is the poset consisting of an antichain on $r$ elements with a largest element adjoined. The factorial functions for $\mathbf{C}_{r}$ are $B(n)=n!$ and $C(n)=r^{n-1}(n-1)!$.
  • b.
    The affine geometry $\mathbb{A}(q)$, ordered by inclusion, may be defined as
    $\mathbb{A}(q)=\{U\setminus H:U\in\mathbb{B}(q)\},$
    where $H$ is the subspace of $V(q)$ spanned by $\{e_{2},e_{3},\ldots\}$. For positive integers $r$, the rank $r$ elements of $\mathbb{A}(q)$ are precisely the non-empty sets of the form $U\setminus H$, where $U$ is an element of $\mathbb{B}(q)$ of rank $r$, see [Oxl11, Chapter 6.2]. It follows that $\mathbb{A}(q)$ is Sheffer (see [Oxl11, Proposition 6.2.5]), and that the factorial functions are
    $B(n)=\mathbf{(n)}!\quad\mbox{ and }\quad C(n)=q^{n-1}\cdot\mathbf{(n-1)}!.$
    Indeed, the maximal chains $\varnothing<x_{1}<\cdots<x_{n}$ in $\mathbb{A}(q)$ are in one-to-one correspondence with the maximal chains $(0)<x_{1}<\cdots<x_{n}$ in $\mathbb{B}(q)$ for which $x_{1}\not\subseteq H$. Hence we should choose $x_{1}$ in $\mathbf{\binom{n}{1}}-\mathbf{\binom{n-1}{1}}=q^{n-1}$ ways, and then the chain $x_{1}<x_{2}<\cdots<x_{n}$ in $B(n-1)=\mathbf{(n-1)}!$ ways.
If $P$ is a Sheffer poset, then
$r_{n,k}(P)=\frac{C(n)}{C(k)\cdot B(n-k)},\quad\quad 0<k\leq n,$
see [ER95].
Theorem 8.8.
Let $P$ be a Sheffer poset with factorial functions $B$ and $C$, and let
$b(z)=\sum_{n=0}^{\infty}\frac{z^{n}}{B(n)}\quad\mbox{ and }\quad c(z)=\sum_{n=0}^{\infty}\frac{z^{n}}{C(n)}.$
Then
$\sum_{n=0}^{\infty}d_{n}(t)\frac{z^{n}}{C(n)}$ $=1+\frac{1-t}{b(tz)-tb(z)}-\frac{c(tz)-tc(z)}{b(tz)-tb(z)},$
$\sum_{n=0}^{\infty}H_{n}(t)\frac{z^{n}}{C(n)}$ $=\frac{(1-t)b(z)}{b(tz)-tb(z)}+\frac{c(z)\cdot b(tz)-c(tz)\cdot b(z)}{b(tz)-tb(z)},$
$\sum_{n=0}^{\infty}A_{n}(t)\frac{z^{n}}{C(n)}$ $=1+t\frac{c(z)-c(tz)}{b(tz)-tb(z)},$
$\sum_{n=0}^{\infty}G_{n}(t)\frac{z^{n}}{C(n)}$ $=\frac{c(z)\cdot b(tz)-tb(tz)\cdot b(z)}{b(tz)-tb(z)}.$
Proof.
We prove the identity for $d_{n}$, the others follows similarly. Let $R^{\prime}=(r^{\prime}_{n,k})_{n,k=0}^{\infty}=(r_{n,k}\cdot C(k)/C(n))_{n,k=0}^{\infty}$, i.e.,
$r^{\prime}_{n,k}=\begin{cases}1/C(n),&\mbox{ if }k=0,\\ 1/B(n-k),&\mbox{ if }0<k\leq n,\\ 0,&\mbox{ otherwise. }\end{cases}$
By Proposition 3.3, $d_{n}/C(n)=d_{n}[R^{\prime}]$. From (3.1), we derive
$t^{n}/C(n)-t^{n}/B(n)+\sum_{k=0}^{n}d_{k}\frac{t^{n-k}}{B(n-k)}=t/C(n)-t/B(n)+t\sum_{k=0}^{n}d_{k}\frac{1}{B(n-k)},\quad n\geq 0,$
from which the proposed identity follows after some manipulations. ∎
The next corollaries follow from Theorem 8.8 and Example 8.7.
Corollary 8.9.
Let $\{H_{n}\}_{n=0}^{\infty}$ be the Chow polynomials for the $r$-cubical lattice, $r>0$. Then
$\sum_{n\geq 0}H_{n+1}(t)\frac{z^{n}}{r^{n}n!}=\frac{e^{z/r+tz}-te^{tz/r+z}}{e^{tz}-te^{z}}.$
Corollary 8.10.
Let $\{H_{n}\}_{n=0}^{\infty}$ be the Chow polynomials for the affine geometry $\mathbb{A}(q)$. Then
$\sum_{n\geq 0}H_{n+1}(t)\frac{z^{n}}{q^{n}\mathbf{(n)}!}=\frac{e_{q}(z/q)\cdot e_{q}(tz)-te_{q}(tz/q)\cdot e_{q}(z)}{e_{q}(tz)-te_{q}(z)}.$
@@ 8.2. Symmetric and supersymmetric functions
Let $\mu\subseteq\lambda$ be integer partitions. Then the supersymmetric skew Schur function, $s_{\lambda/\mu}(\mathbf{x}/\mathbf{y})$, may be defined by the Jacobi-Trudi identity as
@@
We say that $f(\mathbf{x}/\mathbf{y};t)$ is super Schur $\gamma$-positive if each $\gamma_{k}(\mathbf{x}/\mathbf{y})$ has a nonnegative expansion in super Schur functions. Notice that $f(\mathbf{x}/\mathbf{y};t)\in\Lambda(\mathbf{x}/\mathbf{y})$ is super Schur $\gamma$-positive if and only if $f(\mathbf{x}/0;t)$ is Schur $\gamma$-positive if and only if $f(0/\mathbf{y};t)$ is Schur $\gamma$-positive. This is because the maps from supersymmetric functions to symmetric functions defined by
@@
$s_{\lambda/\mu}(\mathbf{x}/\mathbf{y})=\sum_{\nu}c_{\mu\nu}^{\lambda}\cdot s_{\nu}(\mathbf{x}/\mathbf{y}),$